Ticket Number: HEA-NOC/20040115-6 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20040115 13:15 UTC Problem Start: 20040115 10:00 UTC
Ticket Closed: Problem End:
Packet loss noted to LYIT.
Intermitent loss of connectivity to/from LYIT.
20040115 13:12 BN Attempted to contact John O'Raw, but no answer so left
message. Connection has looked better over the last hour. Will attempt
to recontact after lunch if he does not return call.
20040116 09:39 BN LYIT confirmed in the middle of a major outage. ITnet
are looking for interim solutions while problem is fixed. DCU, as MX
backup, have been informed of increased mail volumes. Outage is expected
to continue into the weekend.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20040115-6